hello. i got 85 marks in jee advanced. i hav sc quota. can i get into iit ?

hello.
i got 85 marks in jee advanced.
i hav sc quota.
can i get into iit ?

Grade:12

1 Answers

Sameer Sharma
6 years ago
sure Keep it up U will never get into it.
Hello student, please check the solution of your question. Given below:
cosA+cosB=1
2cos(A+B/2)cos(A-B/2)=1
2(root3/2)cos(A-B/2)=1
cos(A-B/2)=1/root3
cos(A-B)=2cos2(A-B/2)-1= 2*1/3-1= -1/3